# Turning English-centric LLMs Into Polyglots: How Much Multilinguality Is Needed?

### v1

@@ Acknowledgements
We sincerely thank our friends and colleagues, Fabian Aiolfi, Thea Bächler, Anastassia Shaitarova, Finnur Ágúst Ingimundarson, Cui Ding, Andrianos Michail, Janine Aeberhard, Arnisa Fazla, Farhad Nooralahzadeh, Omnia Ibrahim, Xuan Lam and Manasi Muglikar, for helping us with language-specific questions and validating translations used in our experiments, as well as Lena Bolliger and Patrick Haller for providing valuable feedback. Rico Sennrich acknowledges funding by the Swiss National Science Foundation (project MUTAMUR; no. 213976).
@@ Appendix A Pretraining Data for English-centric LLMs
LLMs are pretrained on enormous amounts of unlabelled text gathered from existing corpora and the web. With the notable exception of BLOOM (Le Scao et al., 2023), very limited information is offered about the distribution of languages represented in datasets used to train LLMs and the subsequent performance on non-English languages. Table 1 provides an overview of the document-level language distributions of the LLMs used in this paper. For Llama 2, the information is taken from the original paper Touvron et al. (2023b), in which the authors analyse the training data using a fastText language classifier on corpus documents with a threshold of 0.5. For GPT-3, we use the official dataset statistics made available on Github [^12] https://github.com/openai/gpt-3/blob/master/dataset_statistics/languages_by_document_count.csv which provide document-level language identification information. To gather statistics for Falcon, we inspected a sample of the RefinedWeb dataset Penedo et al. (2023a) that was constructed to train this family of models. [^13] https://huggingface.co/datasets/tiiuae/falcon-refinedweb Using the OpenLID fastText model from Burchell et al. (2023), we identify the most frequent document-level languages based on approximately 320 million examples from this corpus. Languages are identified with a confidence threshold of $\geq$ 0.6 and predictions below this threshold are aggregated under ‘unknown’.
@@
1 English 89.70% 1 English 94.790% 1 English 93.69%
2 unknown 8.38% 2 unknown 4.944% 2 German 1.20%
3 German 0.17% 3 Standard Malay 0.031% 3 French 1.02%
4 French 0.16% 4 Swahili 0.027% 4 Portuguese 0.64%
5 Swedish 0.15% 5 French 0.016% 5 Italian 0.58%
6 Chinese 0.13% 6 Portuguese 0.012% 6 Spanish 0.51%
7 Spanish 0.13% 7 Chinese (Traditional) 0.011% 7 Dutch 0.37%
8 Russian 0.13% 8 Sardinian 0.011% 8 Polish 0.25%
9 Dutch 0.12% 9 Dutch 0.009% 9 Japanese 0.25%
10 Italian 0.11% 10 Spanish 0.009% 10 Danish 0.16%
11 Japanese 0.10% 11 Italian 0.008% 11 Norwegian 0.15%
12 Polish 0.09% 12 German 0.008% 12 Romanian 0.13%
13 Portuguese 0.09% 13 Danish 0.005% 13 Finnish 0.13%
14 Vietnamese 0.08% 14 Russian 0.005% 14 Chinese (Simplified) 0.12%
15 Ukranian 0.07% 15 Indonesian 0.005% 15 Russian 0.11%
16 Korean 0.06% 16 Swedish 0.005% 16 Czech 0.10%
17 Catalan 0.04% 17 Somali 0.005% 17 Swedish 0.06%
18 Serbian 0.04% 18 Vietnamese 0.004% 18 Hungarian 0.06%
19 Bahasa Indonesian 0.03% 19 Venetian 0.004% 19 Chinese (Traditional) 0.04%
20 Czech 0.03% 20 Chinese (Simplified) 0.004% 20 Bahasa Indonesian 0.04%
21 Finnish 0.03% 21 Tagalog 0.004% 21 Croation 0.04%
22 Hungarian 0.03% 22 Northern Uzbek 0.003% 22 Turkish 0.04%
23 Norwegian 0.03% 23 Limburgish 0.003% 23 Catalan 0.03%
24 Romanian 0.03% 24 Polish 0.003% 24 Vietnamese 0.03%
25 Bulgarian 0.02% 25 Turkish 0.003% 25 Slovenian 0.03%
26 Danish 0.02% 26 Norwegian Nynorsk 0.003% 26 Estonian 0.02%
27 Slovenian 0.01% 27 Norwegian Bokmal 0.003% 27 Slovak 0.02%
28 Croatian 0.01% 28 Romanian 0.002% 28 Korean 0.02%
n/a Icelandic n/a 30 Greek 0.002% 31 Greek 0.017%
n/a Hindi n/a 60 Bulgarian 0.001% 45 Hindi 0.004%
n/a Greek n/a 72 Icelandic 0.001% 47 Icelandic 0.004%
@@
Table 1: Distribution of document languages in pretraining datasets for popular English-centric LLMs. For Llama 2 and GPT-3, statistics are taken from the original works. For Falcon, statistics are estimated based on a sample of the RefinedWeb corpus (roughly 320 million documents). Note that for Llama 2 and Falcon, ‘unknown’ indicates texts identified below a predefined confidence threshold (0.5 for Llama 2, 0.65 for our analysis with RefinedWeb). Touvron et al. (2023b) state that this data partially pertain to programming code. Target languages for our experiments are highlighted in bold.
Appendix B Translating AlpacaEval Prompts
@@
Figure 8: Prompt template used to translate AlpacaEval with GPT-3.5-Turbo. Values in curly braces represent placeholders. The value specified for ‘target language’ is English name of the language (e.g., ‘German’) we are translating into, except for Greek and Chinese, where we specify standard modern Greek and Mandarin Chinese respectively.
@@
To translate AlpacaEval prompts from English into each target language, we query GPT-3.5-Turbo using the template in Figure 8. To validate the translated prompts, we manually inspected a sample of the outputs in various languages. This inspection revealed that the translations were typically decent, although often included literal translations for metaphorical expressions rather than how a native speaker might express themselves. For instance, English ‘bullet points’ was translated literally into Hindi, rather than an arguably more appropriate phrasing such as ‘important points’. Additionally, in languages that distinguish between formal and informal or gendered pronouns (e.g., German, French, Hindi), the formal and male forms are dominant, which may be less representative of how native speakers actually interact with LLM chatbots.
Appendix C Direct vs. Translated Evaluation for Non-English Chat Responses
@@
While using a powerful LLM to evaluate the outputs of other models has been shown to achieve reasonable agreement with human judgements in English (Zheng et al., 2023; Chiang and Lee, 2023), it is unclear whether this agreement transfers to all languages under investigation. Recent work by Hada et al. (2023) has shown that agreement between human and LLM judges tends to be lower for non-English languages, especially in the case of low-resource and non-Latin scripted languages, where the LLM judge tends to be overly optimistic in its assessment. However, for certain assessment criteria, such as linguistic acceptability and general content quality, they also confirm that inter-annotator agreement between LLM-based evaluators and humans is inline with that of multiple human annotators.
@@
To investigate this potential bias, we compared scores assigned by the LLM judge on model-generated responses directly in each non-English target language and on their English translations. For each non-English prompt-response pair, we use GPT-3.5-Turbo (gpt-3.5-turbo-1106) and the prompt shown in Figure 8, specifying English as the target language. The resulting translated response is then paired with its corresponding AlpacaEval prompt in its original English form and presented to the LLM judge using the prompt in Figure 16.
@@
Figure 9 shows that the distribution of assigned scores in the direct and translated evaluation settings is very similar for most languages. For languages that use non-Latin scripts (e.g., Chinese, Russian, Bulgarian), we observe that GPT-3.5-Turbo tends to assign slightly higher scores more frequently when evaluating directly on the non-English prompt/response pairs. This finding agrees with those from Hada et al. (2023) and indicates that LLM-based evaluations in non-English languages tend to be slightly overly optimistic and should be considered with caution. Nevertheless, we observe that the discrepancy between direct and translated evaluation is relatively minor for the languages considered and leads to negligible differences in the overall average helpfulness score. Because of this, and to keep evaluation costs to a minimum, we opt to use the direct evaluation strategy for our experiments.
Appendix D Translating Non-English Guanaco Training Examples into English
You are a helpful assistant. Translate the following conversation between a human and an AI assistant into {{target_language}}. Keep the structure of the original text and preserve things like code, names and role labels (e.g. <S1>, <S2>). Please ensure that your response contains only the translated text. The translation must convey the same meaning as the original and be natural for native speakers with correct grammar and proper word choices. Your translation must also use exact terminology to provide accurate information even for the experts in the related fields. Original: {{training_instance}} Translation into {{target_language}}
Figure 10: Prompt template used to translate Guanaco training instances from English into various target languages.
In order to investigate the effect of language diversity compared to instruction diversity, we translate a subset of Guanaco’s English training examples into non-English target languages and use these to create MT-based Multi-$i$ instruction-tuning datasets. By default, speaker roles in Guanaco are denoted with ‘### Human:’ and ‘### Assistant’. To ensure that these are never translated and the dialogue structure is maintained, we substitute them with special tokens ‘<S1>’ and ‘<S2>’ and explicitly tell the model to leave these tokens intact (see Figure 10). Before training, the special tokens back to their original form.
Appendix E Results with Falcon 7b
@@ E.1 Chat
Figure 11 shows the helpfulness scores assigned by our LLM judge for Falcon 7b given incremental multilingual instruction tuning across all target languages. Similarly to our results with Llama 2 7b (see Figure 2), cross-lingual transfer is elicited after finetuning with relatively few languages, and no additional gains observed when including more than three languages. That said, Falcon 7b appears to show strong performance on French, even without multilingual finetuning, indicating that, despite being an English-centric model, it has strong capabilities in French out of the box. For Spanish, German and Chinese, performance is comparable to that of Llama 2 7b. However, for all other languages, responses are often ranked least helpful, indicating that Falcon 7b’s multilingual capabilities are limited strictly to major European languages and Chinese.
E.2 Extractive Question Answering
@@
Figure 12 shows the results of Falcon 7b on XQuAD. While performance is generally lower than that of Llama 2 7b on this task (see Figure 3), we observe a similar, albeit weaker, effect of multilingual finetuning within the supported languages (Spanish, German, Chinese).
E.3 Commonsense Reasoning
@@
Figure 13 shows the results of Falcon 7b on X-CSQA. Strikingly, in contrast to the results achieved with Llama 2 7b (Figure 4), Falcon 7b fails to score above random performance across all target languages. Regarding the effect of multilingual instruction tuning, we again see that it fails to deliver any performance improvements on this highly structured task.
E.4 XNLI

### v2

@@ Acknowledgements
We sincerely thank our friends and colleagues, Fabian Aiolfi, Thea Bächler, Anastassia Shaitarova, Finnur Ágúst Ingimundarson, Cui Ding, Andrianos Michail, Janine Aeberhard, Arnisa Fazla, Farhad Nooralahzadeh, Omnia Ibrahim, Xuan Lam, Manasi Muglikar, Anne Göhring, and Alessia Battisti for helping us with language-specific questions and validating translations used in our experiments, as well as Lena Sophia Bolliger and Patrick Haller for providing valuable feedback on earlier versions of this paper. Rico Sennrich acknowledges funding by the Swiss National Science Foundation (project MUTAMUR; no. 213976).
@@ Appendix A Pretraining Data for English-centric LLMs
With the notable exception of BLOOM Le Scao et al. (2023) and the recent OLMo models Groeneveld et al. (2024), very limited information is offered about the distribution of languages represented in datasets used to train LLMs and the subsequent performance on non-English languages. Table 1 provides an overview of the document-level language distributions of the LLMs used in this paper. For Llama 2, the information is taken from the original paper Touvron et al. (2023), in which the authors analyse the training data using a fastText language classifier on corpus documents with a threshold of 0.5. For GPT-3, we use the official dataset statistics made available on GitHub which provide document-level language identification information. [^16] https://github.com/openai/gpt-3/blob/master/dataset_statistics/languages_by_document_count.csv To gather statistics for Falcon, we inspected a sample of the RefinedWeb dataset Penedo et al. (2023) that was constructed to train this family of models. [^17] https://huggingface.co/datasets/tiiuae/falcon-refinedweb Using the OpenLID fastText model from Burchell et al. (2023), we identify the most frequent document-level languages based on approximately 320 million examples from this corpus. Following Touvron et al. (2023), languages are identified based on a confidence threshold of 0.5 and predictions below this threshold are aggregated under ‘unknown’. Note, that while this language distribution has a higher concentration of English data than both Llama 2 and GPT-3, Almazrouei et al. (2023) combine the RefinedWeb with additional curated corpora to train the Falcon models, and thus the true language distributions for this model may differ from these estimates.
@@
1 English 89.70% 1 English 95.751% 1 English 93.69%
2 unknown 8.38% 2 unknown 3.902% 2 German 1.20%
3 German 0.17% 3 Standard Malay 0.045% 3 French 1.02%
4 French 0.16% 4 Swahili 0.038% 4 Portuguese 0.64%
5 Swedish 0.15% 5 French 0.019% 5 Italian 0.58%
6 Chinese 0.13% 6 Sardinian 0.014% 6 Spanish 0.51%
7 Spanish 0.13% 7 Portuguese 0.014% 7 Dutch 0.37%
8 Russian 0.13% 8 Dutch 0.01% 8 Polish 0.25%
9 Dutch 0.12% 9 Chinese (Traditional) 0.013% 9 Japanese 0.25%
10 Italian 0.11% 10 Spanish 0.010% 10 Danish 0.16%
11 Japanese 0.10% 11 Italian 0.010% 11 Norwegian 0.15%
12 Polish 0.09% 12 German 0.009% 12 Romanian 0.13%
13 Portuguese 0.09% 13 Danish 0.008% 13 Finnish 0.13%
14 Vietnamese 0.08% 14 Indonesian 0.007% 14 Chinese (Simplified) 0.12%
15 Ukranian 0.07% 15 Somali 0.006% 15 Russian 0.11%
16 Korean 0.06% 16 Swedish 0.006% 16 Czech 0.10%
17 Catalan 0.04% 17 Russian 0.006% 17 Swedish 0.06%
18 Serbian 0.04% 18 Venetian 0.005% 18 Hungarian 0.06%
19 Bahasa Indonesian 0.03% 19 Vietnamese 0.005% 19 Chinese (Traditional) 0.04%
20 Czech 0.03% 20 Northern Uzbek 0.005% 20 Bahasa Indonesian 0.04%
21 Finnish 0.03% 21 Limburgish 0.005% 21 Croation 0.04%
22 Hungarian 0.03% 22 Tagalog 0.005% 22 Turkish 0.04%
23 Norwegian 0.03% 23 Chinese (Simplified) 0.005% 23 Catalan 0.03%
24 Romanian 0.03%
24 Vietnamese 0.03%
25 Bulgarian 0.02% 27 Norwegian Bokmal 0.003% 25 Slovenian 0.03%
26 Danish 0.02%
26 Estonian 0.02%
27 Slovenian 0.01% 33 Finnish 0.003% 27 Slovak 0.02%
28 Croatian 0.01%
28 Korean 0.02%
35 Greek 0.002%
n/a Greek n/a
31 Greek 0.017%
46 Hungarian 0.002%
n/a Hindi n/a
45 Hindi 0.004%
51 Slovenian 0.001%
n/a Icelandic n/a
47 Icelandic 0.004%
68 Bulgarian 0.001%
@@
Table 1: Distribution of document languages in pretraining datasets for popular English-centric LLMs. For Llama 2 and GPT-3, statistics are taken from the original works. For Falcon, statistics are estimated based on a sample of the RefinedWeb corpus (roughly 320 million documents). Note that for Llama 2 and Falcon, ‘unknown’ indicates texts identified below a predefined confidence threshold of 0.5. Touvron et al. (2023) state that this data partially pertains to programming code. Target languages considered in our experiments are highlighted in bold.
Appendix B Translating Training and Evaluation Data
As part of our experiments and evaluation, we generate translations of existing datasets. All original datasets are licensed under the Apache License 2.0. Since translations are generated with GPT-3.5-Turbo, OpenAI’s usage policy applies to the resulting versions, which we make available for future research.
B.1 AlpacaEval Prompts
@@
Figure 9: Prompt template used to translate AlpacaEval with GPT-3.5-Turbo. Values in curly braces represent placeholders. The value specified for ‘target language’ is the English name of the language (e.g., ‘German’) we are translating into, except for Greek and Chinese, where we explicitly specify standard modern Greek and Mandarin Chinese, respectively.
@@
For the evaluation of single-turn dialogue in non-English languages, we translate AlpacaEval prompts from English into each of our considered target languages using GPT-3.5-Turbo and the template in Figure 9. Manual inspection of a sampled subset of these translated prompts in various languages revealed that the translations were typically decent, although often included literal translations for metaphorical expressions rather than how a native speaker might express themselves. For instance, English ‘bullet points’ was translated literally into Hindi, rather than an arguably more appropriate phrasing such as ‘important points’. Additionally, in languages that distinguish between formal, informal, or gendered pronouns (e.g., German, French, etc.), the formal and male forms are dominant. While these characteristics may be not truly representative of how native speakers actually interact with LLMs, we consider these to be potentially valid queries for general-purpose chat models.
B.2 Guanaco Training Examples
You are a helpful assistant. Translate the following conversation between a human and an AI assistant into {{target_language}}. Keep the structure of the original text and preserve things like code, names and role labels (e.g. <S1>, <S2>). Please ensure that your response contains only the translated text. The translation must convey the same meaning as the original and be natural for native speakers with correct grammar and proper word choices. Your translation must also use exact terminology to provide accurate information even for the experts in the related fields. Original: {{training_instance}} Translation into {{target_language}}:
Figure 10: Prompt template used to translate Guanaco training instances from English into various target languages to investigate the effect of language diversity vs. instruction diversity.
To investigate the effect of language diversity compared to instruction diversity (§5.1), we translate a subset of Guanaco’s English training examples into non-English target languages and use these to create MT-based Multi-$i$ instruction-tuning datasets. By default, speaker roles in Guanaco are denoted with ‘### Human:’ and ‘### Assistant’. To ensure that these are never translated and the dialogue structure is maintained, we substitute them with special tokens ‘<S1>’ and ‘<S2>’ and explicitly tell the model to leave these tokens intact (see Figure 10). Before training, we map the special placeholder tokens back to their original form.
Appendix C Hardware Requirements for Instruction Tuning
As mentioned in §3, we use LoRA Hu et al. (2021) for parameter-efficient finetuning. As hyperparameters, we set $R=64$, $\alpha=16$. This results in the number of trainable parameters being roughly 2.3% of the original model size for Llama 2 7B, 1.2% for Llama 2 70B and 1.8% for Falcon 7B. For 7-billion parameter models, we use two NVIDIA GeForce RTX 3090 with 24GB of memory. The time required for each training run is approximately 8 hours. For the larger 70-billion parameter model, discussed in §5.2, we use the same hyperparameters and train on four NVIDIA A100 GPUs with 80GB of memory each. Here, a single training run takes approximately 20 hours.
Appendix D MultiSim Datasets
MultiSim is composed of 34 distinct text simplification data sets aligned at either the document- or sentence-level Ryan et al. (2023), covering 13 different languages. Due to restrictive licensing on some datasets, only a subset are openly accessible. An overview of these datasets is provided in Table 2. As can be seen, domain coverage and the amount of available data vary considerably across languages. To balance out the number of evaluation instances between the different languages, we randomly sample up to 1,371 complex-simple sentence pairs from all available corpora for each language, except for Slovenian, where we use all 939 available instances. While we prioritise sampling from the test and validation splits shared by Ryan et al. (2023), we also draw samples from the designated train splits for Portuguese, German, and Slovenian in order to have a sufficiently large evaluation set for our experiments.
Language Source corpora # instances avail. (test / val / train) # eval. instances
en ASSET Alva-Manchego et al. (2020) 359 / 100 / 19,000 1,371
WikiAuto Jiang et al. (2020) 5,002 / 4,988 / 576,126
ru RSSE Sakhovskiy et al. (2021) 1,083 / 97 / 3,182 1,371
RuAdapt Encyclopedia Dmitrieva et al. (2021) 839 / 840 / 7,782
RuAdapt Fairytale Dmitrieva et al. (2021) 31 / 31 / 248
RuWikiLarge Sakhovskiy et al. (2021) 312 / 678 / 246,978
de GEOLino Mallinson et al. (2020) 81 / 82 / 958 1,371
TextComplexityDE Naderi et al. (2019) 26 / 28 / 208
fr CLEAR Grabar and Cardon (2018) 100 / 294 / 4,196 1,371
WikiLarge FR Cardon and Grabar (2020) 345 / 878 / 296,402
pt PorSimples Aluísio and Gasperin (2010) 420 / 420 / 6,290 1,371
it AdminIT Miliani et al. (2022) 49 / 48 / 588 1,371
SIMPITIKI Wiki Tonelli et al. (2016) 160 / 146 / 1,436
PaCCSS-IT Brunato et al. (2016) 1,061 / 1,061 / 60,485
Teacher Brunato et al. (2015) 17 / 17 / 136
Terence Brunato et al. (2015) 101 / 102 / 809
sl SloTS Gorenc and Robnik-Šikonja (2022) 96 / 94 / 749 939
Table 2: Overview of datasets included in the MultiSim benchmark Ryan et al. (2023). # instances avail. denotes the total number of complex-simple sentence pairs available for each corpus in the MultiSim benchmark. # eval. instances denotes the number of items sampled from all test, validation, and train splits (where necessary) for language used to evaluate performance on sentence simplification.
Appendix E Direct vs. Translated Evaluation for Non-English Dialogue Responses
@@ Appendix D MultiSim Datasets
While using a powerful LLM to evaluate the outputs of other models has been shown to achieve reasonable agreement with human judgements in English Zheng et al. (2023); Chiang and Lee (2023), it is unclear whether this agreement transfers to all languages under investigation. Recent work by Hada et al. (2024) has shown that agreement between human and LLM judges tends to be lower for non-English languages, especially in the case of low-resource and non-Latin scripted languages, where the LLM judge tends to be overly optimistic in its assessment. However, for certain assessment criteria, such as linguistic acceptability and general content quality, they also confirm that inter-annotator agreement between LLM-based evaluators and humans is in line with that of multiple human annotators.
@@
To investigate this potential bias, we compared scores assigned by the LLM judge on model-generated responses directly in each non-English target language and on their English translations. For each non-English prompt-response pair, we translate it into English using GPT-3.5-Turbo (gpt-3.5-turbo-1106) and the prompt shown in Figure 9. We pair the resulting translated responses with their corresponding English AlpacaEval prompts and repeat our LLM judge evaluation.
@@
Figure 11 shows that the distribution of assigned scores in the direct and translated evaluation settings is very similar for most languages. For languages that use non-Latin scripts (e.g., Chinese, Russian, Korean, etc.), we observe that the LLM judge tends to assign slightly higher scores more frequently when evaluating directly on the non-English prompt/response pairs. This finding agrees with those from Hada et al. (2024) and indicates that LLM-based evaluations in non-English languages can be overly optimistic and should be considered with caution. Nevertheless, we observe that the discrepancy between direct and translated evaluation is relatively minor for the languages considered and leads to negligible differences in the overall average helpfulness score. Based on these results, we opt to use the direct evaluation strategy for our experiments, which has the added benefit of avoiding the introduction of potential translation errors and keeping the cost of evaluation to a minimum.
Appendix F Results with Falcon 7b
@@ F.1 Single-turn Dialogue
Figure 12 shows the helpfulness scores assigned by the LLM judge for Falcon 7b given incremental multilingual instruction tuning across all target languages. Similarly to our results with Llama 2 7b (cf. Figure 2), cross-lingual transfer can be elicited after finetuning with relatively few languages, and no additional gains observed when including more than three languages. Interestingly, Falcon 7b appears to show strong performance on French, even without multilingual finetuning, indicating that, despite being an English-centric model, it has strong capabilities in French out of the box. For Spanish, German, and Chinese, performance is comparable to that of Llama 2 7b. However, for all other languages, responses are often ranked least helpful, indicating that Falcon 7b’s multilingual capabilities are limited strictly to major European languages and Chinese.
F.2 Sentence Simplification
Figure 13 depicts Falcon 7b’s performance on the sentence simplification task given multilingual instruction tuning. In general, the results emulate those seen with Llama 2 7b. Performance on English remains uniform as multilinguality increases, while for non-English target languages, the largest jump in performance happens when moving from monolingual to bilingual instruction tuning, and plateauing again with just three languages. The notable exceptions here are French and Italian, for which Falcon 7b appears to provide greater support even in the monolingual instruction tuning setting, echoing the results from the single-turn dialogue task.
F.3 Extractive Question Answering
@@
Figure 14 shows the results of Falcon 7b on XQuAD. While performance is generally lower than that of Llama 2 7b on this task (cf. Figure 4), we observe a similar, albeit weaker, effect of multilingual finetuning within the supported languages (es, de, zh).
F.4 Commonsense Reasoning
@@
Figure 15 shows the results of Falcon 7b on X-CSQA. Strikingly, in contrast to the results achieved with Llama 2 7b (cf. Figure 5), Falcon 7b fails to score above random performance across all target languages. Regarding the effect of multilingual instruction tuning, we again see that it fails to deliver any performance improvements on this highly structured task. Degradation on Russian and Vietnamese, is likely due to the model consistently failing to follow instructions and reproducing part of the question as the response.
F.5 XNLI